国際シンポジウム “Singularity Theory and Commutative Algebra” のお知らせ

９月 11 日 (日) から１６日 (金) まで，科研費基盤 (B) 「多様体の特異点の研究」 (研究代表者；石井志保子) と日本大学文理学部自然科学研究所の共催により標記の国際シンポジウムを開催します．特異点論と可換環論の新しい結びつきの可能性を模索するもので，両分野の交流をますます活発にし，新しい発展を目指すものです．多数のご参加を期待しています．

世話人：石井 志保子, 泊 昌孝, 蔵野 和彦，原 伸生，渡辺 敬一

日程: 2005 年 9 月 11 日 (日） – 16 日（金） (１１日と１６日午後は旅行日です) 場所: 浜名湖カリアック, http://www.curreac.co.jp/

〒 431-1207 静岡県浜松市村櫛町４５９７ 浜名湖頭脳公園内 TEL ０５３−４８４−４１５５, FAX ０５３−４８４−４１５０ (１１日/１６日の午後に東海道線舞阪駅より／へ送迎バスを用意します． )

プログラム

9 月 12 日 （月）

9:00-10:00, Lawrence Ein (Univ. Illinois, Chicago), Contact loci and valuations 10:30-11:30, 高木俊輔 (九州大学), F-thresholds and Bernstein polynomials 11:45-12:45, Tommaso de Fernex (Univ. Michigan),

Singularities of maps and birational rigidity 14:30-15:30, Paul Roberts (Univ. Utah),

Local Cohomology and the Homological Conjectures in Commutative Algebra (I) 16:00-17:00, 蔵野和彦 (明治大学),

Resolution of singularities and the positivity conjecture of Serre 17:15-18:15, Holger Brenner (Univ. Sheffield),

Grothendieck topologies and ideal closure operations

9 月 13 日（火）

9:00–10:00, Jonathan Wahl (Univ. North Carolina),

Universal abelian covers of complex surface singularities (I), 10:30–11:30, 奥間智弘 (群馬工業高専),

Universal abelian covers of certain surface singularities 11:45–12:45, 今野一宏 (大阪大学), On the canonical ring of a surface 14:30–15:30, Lawrence Ein (Univ. Illinois, Chicago),

Asymptotic invariants of line bundles 16:00–17:00, 野間淳 (横浜国大),

Multisecant lines to projective varieties and its application 17:15–18:15, Youngook Choi (KAIST),

Projections and linear syzygies, 19:00–21:00, 懇親会

9 月 14 日 (水)

9:00–10:00, Vasudevan Srinivas (Tata Inst.),

Complete intersection points on affine varieties and zero cycles 10:30–11:30, Jihun Park (POSTECH Pohan),

Q-factoriality of double solids 11:45–12:45, 上原北斗 (京都大学理学研究科),

Affine braid group actions on derived categories of algebraic surfaces 午後 Excursion

9 月 15 日 (木)

9:00–10;00, Paul Roberts (Univ. Utah), Local Cohomology and the Homological Conjectures in Commutative Algebra (II), 10:20–11:20, 伊山修 (兵庫県立大学),

Toward higher dimensional Auslander-Reiten theory 11:30–12:30, 吉野雄二 (岡山大学),

Rigid Cohen-Macaulay modules over a quotient singularity of dimension three 14:00-15:00, Jonathan Wahl (Univ. North Carolina),

Universal abelian covers of complex surface singularities (II), 15:10-16:10, Vijaylaxmi Trivedi (Tata Inst,),

Hilbert-Kunz multiplicities and vector bundles on curves 16:30-17:30, 黒田茂 (京都大学数理研),

Counterexamples to the Fourteenth Problem of Hilbert 17:40-18:40, Guillaume Rond (Univ. Toulouse),

Artin function and diophantine approximation in fields of power series

9 月 16 日 (金)

9:00-10:00, Vasudevan Srinivas (Tata Inst.), Lefschetz theorems for the divisor class group 10:20-11:20, Florin Ambro (京都大学数理研), Finite generation of toric FGA algebras 11:30-12:30, 石井志保子 (東工大), Arcs and valuations

参加ご希望の方は渡辺敬一 (watanabe@math.chs.nihon-u.ac.jp) までお知らせ下さい．